The measure of an angle is 16 less than 3 times the measure of its complement.Find the measure of each angle.
Answer by rm29924(97)

The complement to an angle makes it equal 90 degrees
Let X = the angle's complement
Let 3X-16 = the angle
X + 3X - 16 = 90
combine like terms
4X - 16 = 90
add 16 to both sides
4X = 106
divide both sides by 4 to get the variable X by itself
X = 26.5
check your answer
26.5 + 3(26.5)-16 = 90
26.5 + 79.5 - 16 = 90
90 = 90
the angle's complement = 26.5
3(26.5)-16=63.5
the angle=63.5
answer: 26.5 and 63.5